Wednesday, December 15, 2010

12.15.10

The first 3 of yesterday's stuff went fine. Amazingly I was able to get a license for the driver quickly, so accessing both cameras simultaneously through the same interface is no problem. Current target position seems fine for now, here's an example image with both DMs at the bias voltage.



I'll start looking at performance measures today. Incidentally, here's an illustration of the strange bias WF problem I've been having.



This is a composite of the reference SHWFS image from last night and from this morning. Clearly there's been a change in the focus, which is a mystery since no one's been the lab afaik. This is fixed by adjusting L2 (after the spatial filter), and doing some basic realignment, but its bitch to do correctly.

No comments: